# Robustness of LiDAR-Based Pose Estimation: Evaluating and Improving Odometry and Localization Under Common Point Cloud Corruptions

### v2

@@ Abstract
Accurate and reliable pose estimation, i.e., determining the precise position and orientation of autonomous robots and vehicles, is critical for tasks like navigation and mapping. LiDAR is a widely used sensor for pose estimation, with odometry and localization being two primary tasks. LiDAR odometry estimates the relative motion between consecutive scans, while LiDAR localization aligns real-time scans with a pre-recorded map to obtain a global pose. Although they have different objectives and application scenarios, both rely on point cloud registration as the underlying technique and face shared challenges of data corruption caused by adverse conditions (e.g., rain). While state-of-the-art (SOTA) pose estimation systems achieved high accuracy on clean data, their robustness to corrupted data remains unclear. In this work, we propose a framework to systematically evaluate five SOTA LiDAR pose estimation systems across 18 synthetic real-world point cloud corruptions. Our experiments reveal that odometry systems degrade significantly under specific corruptions, with relative position errors increasing from 0.5% to more than 80%, while localization systems remain highly robust. We further demonstrate that denoising techniques can effectively mitigate the adverse effects of noise-induced corruptions, and re-training learning-based systems with corrupted data significantly enhances the robustness against various corruption types.
@@ I Introduction
Autonomous robots and vehicles rely on real-time pose estimation to navigate, avoid collisions, and perform tasks such as mapping and path planning. LiDAR enables precise environmental perception by generating high-resolution 3D point clouds independent of external lighting conditions. As a result, it is widely used for pose estimation, particularly in scenarios where GNSS signals are unreliable (e.g., urban canyons) or unavailable (e.g., tunnels). Two primary methods are commonly employed for LiDAR-based pose estimation: odometry [1] and localization [2]. LiDAR odometry estimates incremental motion by registering consecutive scans, making it essential for exploring unknown environments. However, odometry suffers from drift due to error accumulation over time. LiDAR localization, on the other hand, aligns online scans with pre-built maps to achieve drift-free global pose estimation. This is particularly crucial in safety-critical applications, such as fully autonomous vehicles that require global localization accuracy at the centimeter level [3]. Despite the different objectives, both tasks share underlying point cloud registration techniques to obtain fine poses.
Point cloud data corruptions can occur in various real-world scenarios, including adverse weather [4], noises from sensor [5] or dusts [6], reflective surfaces [7], and occlusions caused by dynamic environments [8, 9]. Despite SOTA LiDAR-based odometry and localization methods achieving highly accurate results on clean LiDAR datasets, corrupted point cloud data will inevitably bring challenges to their performance [10]. However, the extent to which their performance degrades under these corruptions remains unclear. Severe performance degradation could lead to critical failures, such as robots getting lost [11] and autonomous vehicles driving off-road [12]. Though prior studies explored the robustness [9] and security [13] of LiDAR-based localization systems, there is still a lack of systematic evaluation of how point cloud corruptions impact LiDAR-based pose estimation systems. Our work bridges this gap with three key contributions:
  • A framework to systematically evaluate the robustness of five SOTA LiDAR-based pose estimation systems under 18 common point cloud corruptions.
  • An investigation of mitigation strategies, including denoising and re-training (for learning-based systems).
  • Public release of code and data for reproducibility [^1] https://github.com/boyang9602/LiDARLocRobustness.
@@ II Background and related works
II-A LiDAR-Based Pose Estimation
LiDAR-based pose estimation allows autonomous agents to track their motions (odometry) and get their absolute position and orientation in a map (localization) in 3D environments. The core of LiDAR-based pose estimation is point cloud registration, i.e., aligning LiDAR scans to estimate their 6-degree-of-freedom (DoF) spatial transformations (translation and rotation).
@@ II-A1 Point Cloud Registration
Here, $\mathbf{R}$ is a 3-DoF rotation matrix, and $\mathbf{t}$ is a 3-DoF translation vector. The operation $\mathbf{R}\mathcal{P}+\mathbf{t}$ rotates all points using $\mathbf{R}$ and translates them by $\mathbf{t}$. The function $\text{dist}(\cdot,\cdot)$ quantifies the errors between the corresponding points in $\mathbf{R}\mathcal{P}+\mathbf{t}$ and $\mathcal{T}$. In other words, this equation seeks the optimal pose $\mathbf{R}^{*},\mathbf{t}^{*}$ that aligns the source point cloud $\mathcal{P}$ to the target point cloud $\mathcal{T}$.
Various solutions have been proposed for point cloud registration, evolving from geometry-based methods to learning-based methods. They can be broadly categorized into direct matching, feature-based, and learning-based methods [1]. Direct matching methods, such as Iterative Closest Point (ICP) [14], align raw points by iteratively minimizing the distance between nearest neighbor points. While vanilla ICP struggles with problems such as sensitivity to initialization and noise, modern variants such as KISS-ICP [15] address these limitations and remain competitive currently. Another direct method, Normal Distribution Transform (NDT) [16], partitions the target point cloud into a grid of cells, with each cell modeling the local distribution of points as a Gaussian. Then it finds the optimal pose that transforms the source point cloud to best fit the normal distribution in the target grid. Feature-based methods like LiDAR Odometry and Mapping (LOAM) [17] achieve fast and robust odometry by extracting and matching feature points such as sharp edges and planar surfaces. With the development of deep learning, researchers further improve feature extraction from hand-crafted features to learned features and predict the pose using neural networks [18]. Instead of processing the 3D point cloud directly, [19] projects the points into 2D images and leverages the neural network to extract features from images and predict the pose. The success of Neural Radiance Fields (NeRF) [20] further inspires the recent exploration of implicit neural representations of scenes and advances the LiDAR-based pose estimations [21, 22].
@@ II-A2 LiDAR Odometry
LiDAR odometry tracks the accurate motion of autonomous agents in real-time by aligning consecutive LiDAR scans. Despite the high accuracy of point cloud registration, the long-term trajectory drifts significantly due to accumulated errors. To mitigate the drift, additional algorithms such as loop closure [23] should be applied in post-processing. SOTA odometry systems cover all the approaches mentioned in the previous section [15, 17, 24, 19, 18, 22].
@@ II-A3 LiDAR Localization
LiDAR localization determines the global 6-DoF pose by aligning an online scan with a pre-built map, which is critical for fully autonomous driving systems that require centimeter-level accuracy. While some localization systems estimate only 3-DoF position, we focus on 6-DoF pose estimation in this work. Compared to scan-to-scan registration in odometry, scan-to-map registration in localization is more challenging due to the vast search space. To address this problem, most methods first obtain an initial pose via external signals (e.g., GNSS) [3] or global position recognition [2], then perform the point cloud registration to estimate the fine pose. Although one-stage global pose estimation approaches exist, they generally underperform compared to those with initialization [2]. Similar to odometry, SOTA localization approaches also cover the full spectrum of registration methods [3, 25, 26, 21].
@@ II-A4 Multi-Sensor Fusion Approaches
Multi-sensor fusion (MSF) combines complementary sensors, such as LiDAR, cameras (vision odometry and localization), IMUs (high-frequency motion data), and GNSS (global localization), to enhance pose estimation [3]. In this work, we focus on LiDAR-only pose estimation systems to study LiDAR’s standalone vulnerabilities in pose estimation.
@@ II-B Robustness of LiDAR-based Systems
Previous research has extensively explored the robustness and security of LiDAR-based 3D perception systems, particularly in obstacle detection. Several studies have proposed realistic adverse weather simulations for point cloud, including rain [27, 28], snow [29], and fog [30]. Researchers also systematically benchmarked the LiDAR-based object detection on common point cloud data corruptions [31] and investigated its vulnerability to adversarial attacks [32].
In the domain of LiDAR-based pose estimation, recent works have examined the security of localization and odometry against adversarial attacks [13, 33, 34]. The most relevant study to ours is [9], which proposed a closed-form approximation of localization errors for the ICP algorithm. In contrast, our study systematically evaluates the robustness of multiple pose estimation systems, covering various underlying point cloud registration algorithms, against a comprehensive set of point cloud data corruptions.
@@ III Methodology
This section presents our framework for evaluating and improving the robustness of LiDAR-based pose estimation.
@@
Weather rain, snow, rain + wet ground (rain_wg),
@@
Noise background noise (bg_noise), upsample,
@@
uniform noise in SCS (gau_noise_rad),
uniform noise in SCS (imp_noise_rad),
Density local density increase (local_inc),
@@
TABLE I: The 18 perturbations used by our work from three categories. CCS refers to Cartesian Coordinate System and SCS refers to Spherical Coordinate System. The details will be introduced in Section III.
III-A Evaluating Robustness of LiDAR-Based Pose Estimation
@@
Autonomous robotics and vehicles are deployed in complex and diverse environments which affects the LiDAR performance and, consequently, the accuracy of pose estimation. To rigorously evaluate the robustness of pose estimation systems, our framework employs a broad spectrum of point cloud corruption (18 different types) that simulates adverse weather, internal and external noises, equipment malfunctions, and occlusions. Table I summarizes the list of corruption types.
Weather Perturbations. Adverse weather can negatively affect the quality of LiDAR scans. Randomly distributed droplets in the air from rain, snow, and fog can scatter light and disturb the measurement of LiDAR, and cause corruptions in point cloud. Furthermore, wet ground always co-occurs with rain and snow due to the precipitation, which reflects differently from dry ground. Despite the availability of real-world datasets capturing various adverse weather conditions [35, 36], we opt for simulation methods on clean datasets for the following reasons:
  • Realism. Existing simulation methods are physics model-based and highly resemble real-world phenomena [31, 27, 29, 30].
  • Controllable. In real-world datasets, the severity of adverse weather is typically unknown and cannot be adjusted. Simulation, on the other hand, allows us to parameterize the intensity of these conditions.
  • Sensor limitations. Real-world datasets are limited by their specific sensors, which may be incompatible with certain algorithms [10]. Simulation allows us to apply adverse weather on any dataset.
  • Completeness. No dataset contains the full range of corruptions. Simulation allows us to apply all corruptions on a single dataset for a comprehensive and fair comparison.
@@
Figure 1: An example of rain and wet ground corruption. The field view and reflection intensities are significantly decayed compared to the clean data. The color bar indicates the intensities of the points.
@@
Our framework simulates the following weather conditions considering both the direct impact of weather conditions on LiDAR (e.g., droplets causing scattering), as well as impacts of wet grounds on LiDAR (e.g., reflections).
@@
  • Rain or Snow. We utilize an off-the-shelf weather simulator, i.e., LISA [27, 28] to simulate rain and snow. LISA utilizes a physics model to simulate the decayed return of laser pulses in rain and snow weather.
  • Fog. We utilize Fog Simulation (FS) [30] to simulate fog. FS models fog by modifying the impulse response changes to reflect the impact of different fog densities on the environment and the surrounding objects.
@@
  • Rain or Snow + Wet Ground (rain/snow_wg). We combine LISA and Wet Ground Model from [29] to produce a more realistic simulation than rain/snow alone. The Wet Ground Model builds an optical model to simulate the reflection on the wet ground based on the thickness of the water. Figure 1 shows an example of this corruption.
@@
Noise Perturbations. Noises occur in LiDAR measurement due to internal (e.g., equipment vibration) or external (e.g., dust) factors [37, 38]. We design perturbations on point cloud to simulate the errors attributed to such factors.
@@
Background (CCS) $P_{new}=P\cup\{p_{i}^{\prime}|j=1,2,...,SEV\}$
$p_{i}^{\prime}=\{x_{i}^{\prime},y_{i}^{\prime},z_{i}^{\prime}\},\left\{\begin{array}[]{l}x_{i}^{\prime}\sim\mathcal{U}(x_{\min},x_{\max})\\ y_{i}^{\prime}\sim\mathcal{U}(y_{\min},y_{\max})\\ z_{i}^{\prime}\sim\mathcal{U}(z_{\min},z_{\max})\\ \end{array}\right.$
Upsample (CCS) $P_{new}=P\cup\{p_{i}+\Delta p|\Delta p\sim\mathcal{U}(-0.1,0.1),$
$p_{i}\in RS(P,SEV)\}$
TABLE II: Noise-induced corruptions. SEV (severity) refers to the significance of corruption, whose definition varies based on the type of corruption. C refers to a constant value, whose concrete value varies in different places. RS refers to a random subset function, which samples a subset from the input based on the given $SEV$. $\Delta p$ refers to a point in CCS, so $+\Delta p$ means we change the three coordinates in CCS. $\Delta r$ refers to the range value, $+\Delta r$ means we only change the range value in SCS.
We show the details of noise perturbations in Table II. One frame of point cloud is annotated as $\mathcal{P}=\{p_{i}|i=1,2,...N\}$, where $N$ is the number of points in the point cloud. We apply three distributions of noises (Gaussian, Uniform, and Impulse) to two coordinate systems (Cartesian and Spherical, CCS and SCS for short) to simulate different noise sources. In CCS, one point is represented by $p=\{x,y,z\}$, i.e., Euclidean distance to each planes, while in SCS, one point is represented by $p=\{r,\phi,\theta\}$, where $r$ is the range, i.e., Euclidean distance from the origin to the point, $\phi$ is the azimuthal angle and $\theta$ is the polar angle. Noises in CCS represent inaccuracies caused by the vibration or rotating of the sensor [39] while SCS can better represent inaccurate measurement of time of flight (ToF) [31]. In addition to re-positioning existing LiDAR points, we synthesize new points to simulate environment noises, such as dust in the air, through background noise and upsample applied to CCS. Background noise adds points randomly in the entire space, while upsample adds points close to randomly selected points.
@@
III-B Improving Robustness of LiDAR-Based Pose Estimation
Re-training. Learning-based systems are data-driven and could learn to fit themselves into a data distribution. We expect these systems to learn how to handle the corrupted data by themselves through re-training, which is a common enhancement and defensive strategy for deep learning-based LiDAR-based algorithms [27, 29, 30].
Denoising. Handcraft algorithms are expected to be tuned and adapted by experts once inefficiencies are identified. Therefore such improvements are achieved case by case and lack of generality. Hence in this work, for non-learning-based pose estimation systems, we focus on denoising the data, which is a more general way. Point cloud denoising removes unwanted noises from LiDAR scans to improve their quality. There are three categories of denoising, i.e., filter, optimization, and deep learning-based, each with its advantages and disadvantages [37]. In this work, we adopted a filter-based method, i.e., the bilateral filter by [40], which is reliable, and significantly faster than the alternatives while preserving the intensity with readily available implementation.

### v3

# Evaluating and Improving the Robustness of LiDAR Odometry and Localization Under Real-World Corruptions

@@ Abstract
LiDAR odometry and localization are two widely used and fundamental applications in robotic and autonomous driving systems. Although state‑of‑the‑art (SOTA) systems achieve high accuracy on clean point clouds, their robustness to corrupted data remains largely unexplored. We present the first comprehensive benchmark to evaluate the robustness of LiDAR pose‑estimation techniques under 18 realistic synthetic corruptions. Our results show that, under these corruptions, odometry position errors escalate from 0.5% to more than 80%, while localization performance stays consistently high. To address this sensitivity, we propose two complementary strategies. First, we design a lightweight detection-and-filter pipeline that classifies the point cloud corruption and applies a corresponding filter (e.g., bilateral filter for noise) to restore the point cloud quality. Our classifier accurately identifies each corruption type, and the filter effectively restores odometry accuracy to near-clean data levels. Second, for learning‑based systems, we show that fine-tuning using the corrupted data substantially improves robustness across all tested corruptions and even boosts performance on clean point clouds on one data sequence.
@@ I Introduction
Reliable pose estimation is critical in robotics and autonomous driving systems (ADS), where even minor errors can lead to catastrophic failures in navigation, planning, and control [1, 2]. LiDAR-based pose estimation is widely adopted for its superior reliability and performance, particularly when Global Navigation Satellite System (GNSS) signals are unreliable (e.g., urban canyons) or unavailable (e.g., indoors).
To this end, extensive research has focused on improving LiDAR pose estimation, i.e., odometry and localization. Particularly, LiDAR odometry estimates an agent’s incremental motion between consecutive LiDAR scans, enabling reliable self-tracking and map building in unknown environments. In contrast, LiDAR localization determines an agent’s absolute pose within a pre-built map, supporting accurate navigation in structured settings such as urban road networks. These methods vary in their registration strategies, including direct scan alignment [3, 4], feature-based matching [5], and, more recently, learning-based approaches [6, 7] to improve generalization across diverse environments.
Despite the success, LiDAR odometry and localization face inherent challenges on robustness caused by non-clean LiDAR data due to real-world corruptions. Prior works reveal that point cloud data (PCD) may be corrupted or noisy, caused by real-world environment challenges, such as adverse weather [8], sensor noises [9] or dusts [10], reflective surfaces [11], and occlusions caused by dynamic environments [12]. This issue is particularly pronounced in learning-based approaches, which often lack transparency and interpretability, making it difficult to understand their decision-making processes and limiting their ability to generalize to scenarios that diverge from the training data [13].
While prior studies have examined specific aspects of robustness, e.g., occlusion in  [12] and security threats via adversarial attacks  [14], these works are limited in scope and do not offer a comprehensive robustness benchmark for LiDAR pose estimation systems. Notably, there is no systematic evaluation of how a wide range of real-world point cloud corruptions affect LiDAR odometry and localization, particularly in learning-based approaches, which are known to be more sensitive to data quality. Furthermore, existing efforts fall short in proposing methods to improve the robustness of LiDAR pose estimation against such corruptions.
In this work, we propose a systematic evaluation framework, RobustLOL, to address the research gap regarding the impact of PCD corruptions on LiDAR odometry and localization, especially learning-based methods which are known to be sensitive to such corruptions. Furthermore, to mitigate these adverse effects, we further introduce and experiment with two complementary strategies to enhance robustness. First, we design a novel and lightweight detection‑and‑filter pipeline and validate its effectiveness on noise‑induced corruptions. Second, for learning‑based methods, we show that fine-tuning on corrupted data significantly enhances robustness. Our key contributions are as follows:
  • We propose RobustLOL, a novel comprehensive robustness benchmark for LiDAR odometry/localization systems.
  • We propose a novel lightweight detection-and-filter method which enables accurate, efficient, generalizable classification and restoration of corrupted PCD.
  • We investigate various mitigation strategies, including detection-and-filter and fine-tuning to defend the LiDAR odometry/localization systems against PCD corruptions.
  • Public release of code and data for reproducibility [^1] https://doi.org/10.5281/zenodo.17128983.
Our framework reveals that SOTA LiDAR odometry is sensitive to at least one type of PCD corruptions. Our detection-and-filter can effectively defend the noise-induced corruptions for non-learning systems and corruption augmented fine-tuning can bring learning-based systems to near-clean level.
@@ II Background and Related Works
II-A LiDAR-Based Odometry and Localization
LiDAR-based odometry and localization allow autonomous agents to track their motions (odometry) and get their absolute position and orientation in a map (localization) in 3D environments. The core of LiDAR-based odometry and localization is point cloud registration, the process of aligning two LiDAR scans to determine their relative poses.
@@ II-A1 Point Cloud Registration
where $\mathbf{R}$ is a 3-DoF rotation matrix and $\mathbf{t}$ is a 3-DoF translation vector. The source point cloud is rotated by $\mathbf{R}$ and translated by $\mathbf{t}$. The function $\text{dist}(\cdot,\cdot)$ quantifies the errors between the corresponding points in $\mathbf{R}\mathcal{P}+\mathbf{t}$ and $\mathcal{T}$. In other words, this equation seeks the optimal pose $\mathbf{R}^{*},\mathbf{t}^{*}$ that aligns the source point cloud $\mathcal{P}$ to the target point cloud $\mathcal{T}$.
Various approaches are proposed to solve this problem, evolving from geometry-based matching methods to modern learning-based methods. They can be broadly categorized into direct matching, feature-based, projection-based, and neural distance field-based methods. Direct methods like iterative closest point (ICP) [15, 4] align raw points by iteratively minimizing the distance between nearest neighbor points. Normal distribution transform (NDT) is another classic direct method [3], which partitions the target point cloud into a grid of cells, with each cell modeling the local distribution of points as a Gaussian. Instead of minimizing the distances between points, it optimizes the likelihood measure that evaluates how well the source point cloud fits the normal distribution in the target grid. A recent breakthrough is LiDAR Odometry and Mapping (LOAM) [5] which advances the point cloud registration by extracting robust features (e.g., edges and planar regions) from point clouds and matching the features. With the development of deep learning, researchers further improved feature extraction from hand-crafted features to learned features [6, 7]. The success of Neural Radiance Fields (NeRF) [16] in modeling implicit 3D scenes has inspired the efforts to explore of the application in the LiDAR-based pose estimation [17, 18].
@@ II-A2 LiDAR Odometry
LiDAR odometry tracks real-time motion by aligning consecutive scans. While point cloud registration offers high accuracy, accumulated errors lead to significant long-term drift. To correct this, post-processing techniques like loop closure [19] are often used. SOTA odometry systems cover the various approaches as discussed above [4, 5, 20, 7, 6, 18].
@@ II-A3 LiDAR Localization
LiDAR localization determines the global 6-DoF pose by aligning an online scan with a pre-built map, enabling centimeter-level accuracy critical for fully autonomous driving systems. While some localization systems estimate only 3-DoF position, our work focus on 6-DoF pose estimation. Compared to scan-to-scan registration in odometry, scan-to-map registration in localization is more challenging due to the vast search space. To address this problem, most methods first obtain an initial pose via external signals (e.g., GNSS) [21] or global position recognition [22], then perform the point cloud registration to estimate the fine pose. Although one-stage global pose estimation approaches exist, they generally underperform compared to those with initialization [22]. Similar to odometry, SOTA localization approaches also cover the full spectrum of registration methods [21, 23, 24, 17].
@@ II-A4 Multi-Sensor Fusion Approaches
Multi-sensor fusion (MSF) combines complementary sensors, such as LiDAR, cameras (vision odometry and localization), IMUs (high-frequency motion data), and GNSS (global localization), to enhance pose estimation [21]. In this work, we focus on LiDAR-only pose estimation systems to study LiDAR’s standalone sensitivity to noises in pose estimation.
@@ II-B Robustness of LiDAR-based Systems
Previous research has extensively explored the robustness and security of LiDAR-based 3D obstacle detection. Several studies have also introduced realistic adverse weather simulations to evaluate the impact of environmental conditions on PCD, including rain [25, 26], snow [27], and fog [28]. Researchers have also systematically benchmarked 3D obstacle detection under common point cloud corruptions [29] and investigated its vulnerability to adversarial attacks [30, 31].
For LiDAR pose estimation, recent work has investigated the security of localization and odometry against adversarial attacks [14, 32, 33]. The study most closely related to ours is Laconte et al. [12], which derives a closed-form approximation of the worst-case localization error for the ICP algorithm [15] as a function of the number of corrupted points. Their theoretical bound offers useful insight into how adversarial outliers can degrade a linearized (vanilla) ICP. However, SOTA odometry and localization systems typically employ more advanced ICP variants or alternative registration approaches, and the idealized bounds in [12] do not fully reflect the degradation under realistic corruptions. Our work complements this research gap by performing a systematic empirical evaluation of a broad set of odometry and localization pipelines, including both learning-based and classical registration methods, under a comprehensive suite of realistic corruption scenarios.
@@ III Methodology
In this section, we introduce a unified framework to evaluate LiDAR-based odometry and localization systems under common data corruptions, and we describe two defense strategies: a detection-and-filter pipeline and model fine-tuning on augmented data.
@@
Weather (5) rain, snow, rain + wet ground (rain_wg),
@@
Noise (8) background noise (bg_noise), upsample,
@@
gaussian noise in SCS (gau_noise_rad),
impulse noise in SCS (imp_noise_rad),
Density (5) local density increase (local_inc),
@@
TABLE I: Our robustness framework consists of 18 real-world LiDAR corruptions. CCS refers to Cartesian Coordinate System and SCS refers to Spherical Coordinate System. Their calculation formulas are listed in Table II and III.
III-A RobustLOL: A Robustness Framework for LiDAR Odometry and Localization
@@
Autonomous robots and vehicles are deployed in complex and diverse environments. The performance of LiDAR may degrade under adverse conditions. To rigorously evaluate the robustness of odometry and localization systems, our framework employs 18 different types of PCD corruption (Table I) that simulate adverse weather, internal and external noises, equipment malfunctions, and occlusions.
Weather Perturbations. Adverse weather can negatively affect the quality of LiDAR scans. Randomly distributed airborne droplets from rain, snow, and fog scatter light, corrupting the PCD captured by LiDAR. Furthermore, precipitation results in wet ground that always co-occur with rain or snow, leading to surface reflections that differ significantly from dry ground. Despite the availability of real-world datasets capturing various adverse weather conditions [34, 35], we opt for simulation using clean datasets for three reasons. First, real datasets does not encompass the full spectrum of point-cloud corruptions, whereas simulation enables us to apply every corruptions on the same dataset, ensuring a comprehensive and fair comparison. Second, real recordings are sensor-specific, which can restrict compatibility with certain odometry or localization algorithms [13]; simulation, by contrast, is sensor-agnostic and can be applied to any dataset. Finally, the severity of weather effects in real recordings is fixed and often undocumented, but simulation lets us precisely parameterize corruption intensity and systematically study its impact, leveraging realistic models of fog, rain, and snow [29, 25, 27, 28].
@@
Figure 1: An example of rain and wet ground corruption. The field view and reflection intensities are significantly decayed compared to the clean data. The numbers above the color bar show the intensities of the points with corresponding colors.
@@
Our framework simulates the following weather conditions considering both the direct impact of weather on LiDAR (e.g., droplets causing scattering), as well as impacts of wet grounds on LiDAR (e.g., reflections).
@@
  • Rain or Snow. We use LISA [25, 26], an off-the-shelf rain and snow simulator which use a physics model to simulate the decayed return of laser pulses during rain and snow.
  • Fog. Following previous work [28], we model fog by altering the impulse response, capturing the effect of varying fog densities on the surrounding objects and environment.
@@
  • Rain or Snow + Wet Ground (rain/snow_wg). We combine LISA and Wet Ground Model (WGM) [27] to produce a more realistic simulation than rain/snow alone. WGM builds an optical model to simulate wet ground reflection based on water thickness. We show an example of this corruption in Figure 1.
@@
Noise Perturbations. Noises occur in LiDAR measurement due to internal (e.g., equipment vibration) or external (e.g., dust or strong light) factors [36, 37]. We design perturbations on point cloud to simulate the errors attributed to such factors.
@@
Background $P_{new}=P\cup\{p_{i}^{\prime}|j=1,2,...,SEV\}$
[…]
Upsample $P_{new}=P\cup\{p_{i}+\Delta p|\Delta p\sim\mathcal{U}(-0.1,0.1),$
(CCS) $p_{i}\in RS(P,SEV)\}$
TABLE II: Noise-induced corruptions. SEV (severity) is the significance of corruption, whose definition varies based on the type of corruption. C is a constant value, whose concrete value varies in different scenarios. RS is a random subset function, which samples a subset from the input based on the given SEVSEV. $\Delta p$ is a point in CCS, so $\Delta p$ means we change the three coordinates in CCS. $\Delta r$ refers to the range value, $\Delta r$ means we only change the range value in SCS.
We show the details of noise perturbations in Table II. One frame of point cloud is annotated as $\mathcal{P}=\{p_{i}|i=1,2,...N\}$, where $N$ is the number of points in the point cloud. We apply three distributions of noises (Gaussian, Uniform, and Impulse) to two coordinate systems (Cartesian and Spherical, CCS and SCS for short) to simulate different noise sources. In CCS, one point is represented by $p=\{x,y,z\}$, i.e., Euclidean distance to each planes, while in SCS, one point is represented by $p=\{r,\phi,\theta\}$, where $r$ is the range, i.e., Euclidean distance from the origin to the point, $\phi$ is the azimuthal angle and $\theta$ is the polar angle. Noises in CCS represent inaccuracies caused by the vibration or rotating of the sensor [38] while SCS can better represent inaccurate measurement of time of flight (ToF) [29]. In addition to re-positioning existing LiDAR points, we synthesize new points to simulate environment noises, such as dust in the air or strong external light, through background noise and upsample applied to CCS. Background noise adds points randomly in the entire space, while upsample adds points close to randomly selected points.
@@
III-B Improving the Robustness of LiDAR Odometry and Localization
We propose and experiment with two types of approaches to improve the robustness of such LiDAR systems.
III-B1 For All LiDAR Odometry and Localization.
We propose a novel detection-and-filter pipeline to improve the robustness of LiDAR odometry and localization systems. This pipeline is algorithm-agnostic and can be widely applied to algorithms of LiDAR pose estimations. Our novel pipeline consists of two stages: (1) corruption detection via a DNN-based classifier, and (2) targeted corruption filtering based on the corruption type.
Stage 1 – Corruption Detection. As odometry and localization are real-time applications, we focus on techniques that are inexpensive in terms of time and memory cost. Hence, we first convert each 3D LiDAR scan into a 2D range image via spherical projection similar to [7], creating a $64\times 1024$ grid, where each pixel stores the depth and intensity of the point in that cell closest to the agent. This representation preserves the spatial structure of the scene while enabling efficient convolutional processing.
Using these range images, we train a lightweight CNN (Figure 2) to classify data corruptions (e.g., rain, noises) in LiDAR data on our augmented dataset by minimizing categorical cross-entropy loss. This loss guides the model to improve classification accuracy across different corruption types through successive training epochs.
The CNN consists of three convolutional layers with increasing feature sizes (32, 64, and 128), each followed by batch normalization and ReLU activation. A final adaptive average pooling layer reduces the output to a 128-dimensional vector to produce class scores, where the class with highest probability is selected as the predicted class for the filtering stage. This lightweight architecture is designed for real-time or embedded deployment, balancing classification accuracy with minimal computational overhead.
Figure 2: Architecture of CNN for corruption classification.
Stage 2 – Corruption Removal. After identifying the corruption type, we apply the corresponding filter to mitigate its effects and restore point-cloud fidelity. While our framework supports any paired detector–filter module (e.g., precipitation or density filters), in this work we concentrate on noise-induced corruptions because other point-cloud-based systems are sensitive to noises [29] and denoising methods are mature and effective in removing noises. Thus, once the classifier flags a scan as noise-corrupted, we pass it to a bilateral filter [40] for denoising. This modular design allows additional filters to be plugged in for other corruption categories.
III-B2 For Learning-Based LiDAR Odometry and Localization.
Learning-based approaches inherently adapt to the characteristics of their training data. Hence, to defend these systems against the data corruptions, we augment the training set with perturbations shown to degrade performance by our evaluation framework and then fine-tune the pretrained models on this augmented dataset. We fine-tune the model using the original training recipe from the released checkpoint until it converges on the augmented dataset. This strategy encourages the neural networks to learn invariant features directly from the data distribution, and has been shown to enhance resilience in LiDAR-based systems [25, 27, 28].